# Robustness of quantum algorithms: Worst-case fidelity bounds and implications for design

### v1

@@ III.2 Worst-case fidelity bound
The proof of Theorem III.1 can be found in Appendix E. Theorem III.1 generalizes Theorem II.1 by allowing us to consider incoherent quantum errors, as long as they can be described via a Lindblad equation. This includes a variety of error types such as depolarization, bit- and phase-flips, amplitude damping, dephasing, and more. This generality comes at the cost of additional conservatism. In particular, the bound (42) explicitly involves the number of qubits $n$, and the dependence on $\delta N$ is linear rather than squared. While this means that the obtained worst-case fidelity bounds are less accurate, the conceptual insights of the theoretical results remain, in particular for comparing the robustness of quantum algorithms and for designing algorithms with specified robustness properties. Furthermore, the first bound (39) in Theorem III.1 is applicable even for non-unitary circuit elements, assuming they are invertible. Computing the worst-case fidelity bound (42) for a given algorithm and error model requires obtaining a bound $\gamma$, compare (41). This can be achieved in full analogy to Section II. Hence, we omit the technical details and refer to Appendix C which equally applies in the setup of the present section. Likewise, the main conceptual discussion (Section II.3) as well as the implications for design (Section II.4) carry over to Theorem III.1.
IV Numerical results
In this section, we apply our theoretical framework to study concrete robustness analysis and design problems. In Section IV.1, we analyze the robustness of composite pulse sequences and design new pulses with superior robustness properties. Further, in Section IV.2, we study the robustness of different transpilations of the quantum Fourier transform. The code to reproduce the numerical experiments is publicly available on GitHub [61].
IV.1 Robustness of composite pulses
@@
In the following, we exploit the flexibility of the proposed framework to analyze composite pulses for both systematic and independent errors. To be precise, we apply Theorem II.1 to compare the robustness of a Pauli-$X$ rotation $R_{X}(\frac{\pi}{4})$ and the corresponding composite pulse sequence (46) by Jones [39] against coherent control errors. We compute $\gamma$ via combinatorial optimization, see Appendix B.3. In case of independent errors, we solve the problem (69) directly and then use the bound (14). In case of systematic errors, we exploit the systematic nature to compute tighter bounds by a) exploiting in the computation that all $\theta_{j}$’s are equal and b) using the inequality (11) instead of (14).
@@
Figure 3: The figures show the infidelity $1-F_{\mathrm{wc}}$ with systematic and independent errors depending on the error level $\delta$ for the Pauli-$X$ rotation $R_{X}(\frac{\pi}{4})$, the pulses (46) by Jones [39], and our newly designed pulses (47). The displayed worst-case fidelity bounds are computed based on Theorem II.1.
Finally, we leverage our theoretical analysis tools for designing new composite pulses with superior robustness properties, compare Section II.4. To this end, we optimize over the angles of a $5$-gate composite pulse sequence using Matlab with fmincon, based on the values from (46) as initialization. For a fixed noise level $\delta=0.05$, we maximize the worst-case fidelity for independent coherent control errors. In the optimization, we impose as a constraint that the worst-case fidelity for systematic coherent control errors remains above $99.9995\%$, which is close to the fidelity of the composite pulses from [39]. The resulting sequence is given by
@@
IV.2 Quantum circuit transpilation
@@
Figure 4: Infidelities of the three-qubit QFT circuit as a function of the noise level $\delta$ for different gate sets. Square markers indicate worst-case fidelity bounds computed using Theorem II.1, while circular markers correspond to the worst simulated infidelity obtained from 10,000 randomly drawn noise samples. The number of layers and gates for the transpiled circuits are: A: (34,59), B: (23,42), C: (14,20), D: (9,14).
@@
$H_{\mathrm{e},j}=\theta_{j,1}PII+\theta_{j,2}IPI+\theta_{j,3}IIP\,$, (48)
where $P\in\{X,Y,Z\}$ and $|\theta_{j,k}|\leq\delta/3$ (see Appendix B for details on this choice). Here, we use the notation $ABC\equiv A\otimes B\otimes C$. In the Bloch-sphere picture, these errors correspond to single-qubit rotations about the axis defined by $P$.
@@
Fidelity bounds are computed according to (14), with the parameter $\gamma$ obtained via an optimization procedure (see Appendix B.2). Additional details on the numerical methods used in this section are provided in Appendix F.2. Figure 4 displays the resulting worst-case infidelity bounds for $P=Z$ as a function of the noise level $\delta$ across the four gate sets.
@@
We observe that the robustness against $Z$-errors varies with the choice of gate set. Notably, the number of layers $N$ decreases from circuit A to D, and the ordering of the worst-case fidelity bounds mirrors this trend. To validate the relevance of the bounds, we compare them with the worst-case fidelity obtained by uniformly and independently sampling $\theta_{j}=(\theta_{j,1},\theta_{j,2},\theta_{j,3})\in[-\delta/3,\delta/3]^{3}$ across 10,000 random noise realizations. The observed correspondence between the analytical bounds and the simulation results indicates that the fidelity bound (14) provides a meaningful metric for assessing the relative robustness of different transpiled circuits.
Figure 5: Worst-case infidelity bounds for different Pauli errors (as defined in (48)) across the four gate sets, with $\delta=0.001$. The number of layers and gates for each transpiled circuit (#layers, #gates) are indicated on the x-axis.
We now extend the analysis to include Pauli-$X$ and Pauli-$Y$ errors ($P=X$ and $P=Y$). Figure 5 presents the corresponding worst-case infidelity bounds for all three Pauli error types across the four gate sets. The results demonstrate that the robustness of the QFT circuit is highly sensitive to both the type of error and the choice of underlying gate set. In particular, certain gate sets exhibit greater resilience against specific Pauli errors.
@@
Figure 6: Infidelity bounds for coherent control errors on the transpiled QFT circuits. Theoretical bounds are computed using Theorem II.1 with two different strategies for obtaining $\gamma$: via an optimization problem ($\gamma_{\mathrm{opt}}$, Appendix B.2) and via a norm-based bound ($\gamma_{\mathrm{norm}}$, Appendix B.4). For comparison, the bound from [30, Theorem 2] is also shown. Additionally, we simulate 10,000 randomly sampled noise realizations, reporting both the mean and worst-case infidelity. The noise level is set to $\delta=0.001$. The number of layers and gates of each circuit (#layers, #gates) are indicated on the x-axis.
@@
Next, we analyze the robustness of the transpiled QFT circuits under coherent control errors, as defined in (44). Figure 6 compares different infidelity bounds. We first compute the bound according to Theorem II.1, using two strategies to determine the parameter $\gamma$: First, $\gamma_{\mathrm{opt}}$ obtained via an optimization problem (Appendix B.2), and second, $\gamma_{\mathrm{norm}}$ obtained from a norm-based bound (Appendix B.4). By construction, $\gamma_{\mathrm{opt}}\leq\gamma_{\mathrm{norm}}$, making the infidelity bound with $\gamma_{\mathrm{norm}}$ more conservative. Despite this conservatism, we see that the norm-based bound is sufficient for reliably comparing the robustness of the different circuits against coherent control errors.
@@
For reference, we also include the bound from [30, Theorem 2]. As seen in the figure, Theorem II.1 yields a tighter worst-case bound. A more detailed comparison between Theorem II.1 and [30, Theorem 2] is provided in Appendix D.1.
@@
Finally, we compare these theoretical bounds with simulation results obtained from $10,000$ uniformly sampled noise realizations. While the simulated worst-case fidelities are lower than the theoretical bounds, the relative robustness ranking of the circuits aligns across all considered measures. This indicates that the theoretical fidelity bounds provide a meaningful and quantitatively informative indication of algorithmic robustness under coherent control errors.
V Conclusion
@@
The developed framework opens the door to several promising future research directions. Applying our results to specific design or compilation problems can yield new quantum circuits with superior robustness properties while trading off the influence on the algorithm performance, compare our results in Section IV.1 for first results in this direction. It will also be interesting to explore different error classes in this context such as crosstalk [69] or classes of non-Markovian errors [70]. Finally, open technical challenges include extending the results of Section III to more general (non-invertible) algorithm elements and to improve our theoretical analysis by deriving tighter worst-case fidelity bounds.
@@ Acknowledgments
Appendix A Proof of Theorem II.1
@@
In order to bound the first term on the right-hand side, we use [71, Proposition 9] which states
@@
We now use [52, Inequality (A14)] which states
@@ Appendix B Computation of worst-case fidelity bounds
In the following, we discuss how Theorem II.1 can be implemented computationally to obtain explicit worst-case fidelity bounds for a given quantum algorithm and error model. Throughout this section, we assume that the $H_{\mathrm{e},j}$’s are of the form $H_{\mathrm{e},j}=\sum_{k=1}^{\ell}\theta_{j,k}B_{j,k}$ for some fixed and known matrices $B_{j,k}$ satisfying $\lVert B_{j,k}\rVert\leq 1$ and with unknown but bounded error angles $\theta_{j}=(\theta_{j,1},\dots,\theta_{j,k})$ satisfying $\lVert\theta_{j}\rVert_{\infty}\leq\frac{\delta}{\ell}$, $j=1,\dots,N$. Note that we scale the bounds on $\theta_{j}$ by $\frac{1}{\ell}$ to ensure that (12) is fulfilled. Alternatively, a $2$-norm bound $\lVert\theta_{j}\rVert\leq\delta$ can be considered. The considered setup allows us to model multiple independent errors affecting different qubits: For example, in a $2$-qubit system, independent Pauli-$Z$ rotation errors on each qubit can be captured by $B_{j,1}=Z\otimes I$, $B_{j,2}=I\otimes Z$. Further, we write $G_{j}(\theta_{j})$ and $G(\theta)$ to emphasize the dependence of $G_{j}$ and $G$ on the error $\theta$.
We note that the computational procedures presented in the following can be extended to more general uncertainty descriptions, following ideas from robust quantum control [52]. For example, one can also handle different bounds $\delta$ and numbers $\ell$ for each error Hamiltonian $H_{\mathrm{e},j}$, different vector norms for the bound $\lVert\theta_{j}\rVert\leq\frac{\delta}{\ell}$, or even fully uncertain Hamiltonians $H_{\mathrm{e},j}$ without any structural knowledge of the $B_{j,k}$’s. Moreover, while the main focus of this work lies on independent errors, for which $\theta_{j}$ differs for $j=1,\dots,N$, systematic errors can be analyzed as well, see Section IV.1 for an example.
@@ B.5 Discussion and comparison
It is important to emphasize that none of these approaches scale directly to high numbers of qubits or gates. In particular, each method requires to classically evaluate the matrix $G(\theta)$, whose size grows exponentially in the number of qubits. Moreover, the computation (69) additionally scales exponentially with the number of gates. In the following section, we address this problem via a circuit partitioning strategy, which allows us to compute bounds on $\gamma$ in a scalable fashion, paving the way for relevant problem sizes beyond classical simulation.
@@ D.1 Special case: coherent control errors
In the following, we compare the worst-case fidelity bound from Theorem II.1 to [30], which provides a worst-case fidelity bound for the special case of coherent control errors, compare (44). The main fidelity bound from [30, Theorem 2] states
@@
Note the difference between (15) and (82). When $\delta N$ is small compared to $\gamma$, both infidelity bounds scale as $\delta^{2}N^{2}$. However, for robust algorithms with $\gamma\approx 0$, the bound (15) derived in the present paper scales in a significantly more favorable way with $\delta^{4}N^{4}$. In Section IV.2, we compare both bounds for a concrete set of algorithms, showing that the bound obtained via Theorem II.1 is substantially less conservative than the bound by [30].

### v2

@@ 3.2 Worst-case fidelity bound
The proof of Theorem 3.1 can be found in Appendix E. Theorem 3.1 generalizes Theorem 2.1 by allowing us to consider incoherent quantum errors, as long as they can be described via a Lindblad equation. This includes a variety of error types such as depolarization, bit- and phase-flips, amplitude damping, dephasing, and more. This generality comes at the cost of additional conservatism. In particular, the bound (44) explicitly involves the number of qubits $n$, and the dependence on $\delta N$ is linear rather than squared. While this means that the obtained worst-case fidelity bounds are less accurate, the conceptual insights of the theoretical results remain, in particular for comparing the robustness of quantum algorithms and for designing algorithms with specified robustness properties. Furthermore, the first bound (41) in Theorem 3.1 is applicable even for non-unitary circuit elements, assuming they are invertible. Computing the worst-case fidelity bound (44) for a given algorithm and error model requires obtaining a bound $\gamma$, compare (43). This can be achieved in full analogy to Section 2. Hence, we omit the technical details and refer to Appendix C which equally applies in the setup of the present section. Likewise, the main conceptual discussion (Section 2.3) as well as the implications for design (Section 2.4) carry over to Theorem 3.1.
4 Applications
In this section, we apply our theoretical framework numerically in three robustness analysis and design problems. In Section 4.1, we analyze the robustness of composite pulse sequences and design new pulses with superior robustness properties. Next, in Section 4.2, we study the robustness of different transpilations of the quantum Fourier transform. Finally, in Section 4.3, we study the robustness of a $50$-qubit modular adder circuit by applying circuit partitioning. The code to reproduce the numerical experiments is publicly available on GitHub [60].
4.1 Robustness of composite pulses
@@
In the following, we exploit the flexibility of the proposed framework to analyze composite pulses for both systematic and independent errors. To be precise, we apply Theorem 2.1 to compare the robustness of a Pauli-$X$ rotation $R_{X}(\frac{\pi}{4})$ and the corresponding composite pulse sequence (48) by Jones [42] against coherent control errors. We compute $\gamma$ via combinatorial optimization, see Appendix B.3. In case of independent errors, we solve the problem (72) directly and then use the bound (14). In case of systematic errors, we exploit the systematic nature to compute tighter bounds by a) exploiting in the computation that all $\theta_{j}$’s are equal and b) using the inequality (11) instead of (14).
@@
(a) Systematic errors
(b) Independent errors
Figure 2: The figures show the infidelity $1-F_{\mathrm{wc}}$ with systematic and independent errors depending on the error level $\delta$ for the Pauli-$X$ rotation $R_{X}(\frac{\pi}{4})$, the pulses (48) by Jones [42], and our newly designed pulses (49). The displayed worst-case fidelity bounds are computed based on Theorem 2.1.
Next, we leverage our theoretical analysis tools for designing new composite pulses with superior robustness properties, compare Section 2.4. To this end, we optimize over the angles of a $5$-gate composite pulse sequence using Matlab with fmincon, based on the values from (48) as initialization. For a fixed noise level $\delta=0.05$, we maximize the worst-case fidelity for independent coherent control errors. In the optimization, we impose as a constraint that the worst-case fidelity for systematic coherent control errors remains above $99.9995\%$, which is close to the fidelity of the composite pulses from [42]. The resulting sequence is given by
@@
4.2 Quantum circuit transpilation
@@
Figure 3: Infidelities of the three-qubit QFT circuit as a function of the noise level $\delta$ for different gate sets. Square markers indicate worst-case fidelity bounds computed using Theorem 2.1, while circular markers correspond to the worst simulated infidelity obtained from 10,000 randomly drawn noise samples. The number of gates for the transpiled circuits are: A: 59, B: 42, C: 20, D: 14.
@@
$H_{\mathrm{e},j}=\theta_{j,1}P_{m}$ (50)
for single qubit gates acting on qubit $m$ and
$H_{\mathrm{e},j}=\theta_{j,1}P_{m_{2}}+\theta_{j,1}P_{m_{2}}$ (51)
for two qubit gates acting on qubits $m_{1}$ and $m_{2}$. Here, $P_{m}$ refers to Pauli-error $P\in\{X,Y,Z\}$ acting on qubit $m$. Additionally it holds $|\theta_{j,k}|\leq\delta$ for single qubit gates and $|\theta_{j,k}|\leq\delta$ for two qubit gates (see Appendix B for details on this choice).
@@
Fidelity bounds are computed according to (14), with the parameter $\gamma$ obtained via an optimization procedure (see Appendix B.2). Additional details on the numerical methods used in this section are provided in Appendix F.3.
@@
Figure 3 displays the resulting worst-case infidelity bounds for $P=Z$ as a function of the noise level $\delta$ across the four gate sets. We observe that the robustness against $Z$-errors varies with the choice of gate set. Notably, the number of layers $N$ decreases from circuit A to D, and the ordering of the worst-case fidelity bounds mirrors this trend. To validate the relevance of the bounds, we compare them with the worst-case fidelity obtained by uniformly and independently sampling $\theta_{j,i}$ across 10,000 random noise realizations. The observed correspondence between the analytical bounds and the simulation results indicates that the fidelity bound (14) provides a meaningful metric for assessing the relative robustness of different transpiled circuits.
@@
Figure 4: Infidelity bounds for coherent control errors on the transpiled QFT circuits. Theoretical bounds are computed using Theorem 2.1 with two different strategies for obtaining $\gamma$: via an optimization problem ($\gamma_{\mathrm{opt}}$, Appendix B.2) and via a norm-based bound ($\gamma_{\mathrm{norm}}$, Appendix B.4). For comparison, the bound from [30, Theorem 2] is also shown. Additionally, we simulate 10,000 randomly sampled noise realizations, reporting both the mean and worst-case infidelity. The noise level is set to $\delta=0.001$.
@@
Next, we analyze the robustness of the transpiled QFT circuits under coherent control errors, as defined in (46). Figure 4 compares different infidelity bounds. We first compute the bound according to Theorem 2.1, using two strategies to determine the parameter $\gamma$: First, $\gamma_{\mathrm{opt}}$ obtained via an optimization problem (Appendix B.2), and second, $\gamma_{\mathrm{norm}}$ obtained from a norm-based bound (Appendix B.4). By construction, $\gamma_{\mathrm{opt}}\leq\gamma_{\mathrm{norm}}$, making the infidelity bound with $\gamma_{\mathrm{norm}}$ more conservative. Despite this conservatism, we see that the norm-based bound is sufficient for reliably comparing the robustness of the different circuits against coherent control errors.
@@
For reference, we also include the bound from [30, Theorem 2]. As seen in the figure, Theorem 2.1 yields a tighter worst-case bound. A more detailed comparison between Theorem 2.1 and [30, Theorem 2] is provided in Appendix D.1.
@@
Further, we compare these theoretical bounds with simulation results obtained from $10,000$ uniformly sampled noise realizations. While the simulated worst-case fidelities are lower than the theoretical bounds, the relative robustness ranking of the circuits aligns across all considered measures. This indicates that the theoretical fidelity bounds provide a meaningful and quantitatively informative indication of algorithmic robustness under coherent control errors. Finally, note that the simulated worst-case infidelity of circuit B is slightly larger than that of circuit A although the latter has more gates. The same relation is correctly reflected in the two bounds $\gamma_{\mathrm{opt}}$ and $\gamma_{\mathrm{norm}}$ computed based Theorem 2.1. On the other hand, the existing approach from [30] yields a higher infidelity bound for circuit A, demonstrating the improved accuracy of the present approach.
4.3 Modular adder circuit
One of the most promising application for quantum computing lies in factoring algorithms [68]. A fundamental building block of these algorithms is the modular adder circuit [69], which operates on two $n$-qubit registers to perform the mapping $(a,b)\xrightarrow{}(a,a+b\mod 2^{n})$.
To demonstrate the practical impact of our robustness framework on a relevant large-scale quantum computing task, we apply it to the modular adder architecture proposed in [70] when transpiled into gate set B. We consider a configuration with $n=25$ (totaling 50 qubits), a scale estimated to be sufficient for factoring 2048-bit RSA integers using resources expected to be available in the near future [69].
Since a direct calculation of $\gamma$ is computationally intractable for a system of this size, we partition the circuit into smaller sub-circuits using the methodology described in Appendix C. We set the maximum number of active qubits per partition to three, a value that allows for the efficient numerical calculation of $\gamma_{\mathrm{opt}}$ via the optimization procedure in Appendix B.2. The impact of the partition size on the results is further analyzed in Appendix G.
Figure 5: Infidelities of different Pauli-errors and coherent control errors over the noise level $\delta$ for a 50 qubit modular adder circuit. The infidelity bounds are computed using Theorem 2.1 with $\gamma_{\mathrm{opt}}$ and partitioning into sub-circuits with at most three qubits.
By employing circuit partitioning, we are able to derive a rigorous bound on $\gamma$ even for these large-scale systems. Figure 5 illustrates the resulting infidelity as a function of the noise level $\delta$, determined by Theorem 2.1 for coherent Pauli-type errors (compare equations (50) and (51)) and coherent control errors. Our results indicate that the robustness of the modular adder circuit against Pauli-X and Pauli-Y rotations is slightly improved in comparison to Pauli-Z rotations. Further, it exhibits substantially better robustness against coherent control errors in comparison to Pauli rotations. Thus, implementing the circuit on a hardware platform where Pauli rotation errors (especially Pauli-Z rotations) are less dominant may improve the overall reliability of the circuit. Overall, this numerical example demonstrates that meaningful robustness guarantees can be established even for multi-qubit circuits of practically relevant size.
5 Conclusion
@@
The developed framework opens the door to several promising future research directions. Applying our results to specific design or compilation problems can yield new quantum circuits with superior robustness properties while trading off the influence on the algorithm performance, compare our results in Section 4.1 for first results in this direction. It will also be interesting to explore different error classes in this context such as crosstalk [71] or classes of non-Markovian errors [72]. Finally, open technical challenges include extending the results of Section 3 to more general (non-invertible) algorithm elements and to improve our theoretical analysis by deriving tighter worst-case fidelity bounds.
@@ Acknowledgments
Appendix A Proof of Theorem 2.1
@@
In order to bound the first term on the right-hand side, we use [73, Proposition 9] which states
@@
We now use [37, Inequality (A14)] which states
@@ Appendix B Computation of worst-case fidelity bounds
In the following, we discuss how Theorem 2.1 can be implemented computationally to obtain explicit worst-case fidelity bounds for a given quantum algorithm and error model. Throughout this section, we assume that the $H_{\mathrm{e},j}$’s are of the form $H_{\mathrm{e},j}=\sum_{k=1}^{\ell}\theta_{j,k}B_{j,k}$ for some fixed and known matrices $B_{j,k}$ satisfying $\lVert B_{j,k}\rVert\leq 1$ and with unknown but bounded error angles $\theta_{j}=(\theta_{j,1},\dots,\theta_{j,k})$ satisfying $\lVert\theta_{j}\rVert_{\infty}\leq\frac{\delta}{\ell}$, $j=1,\dots,N$. Note that we scale the bounds on $\theta_{j}$ by $\frac{1}{\ell}$ to ensure that (12) is fulfilled. Alternatively, a $2$-norm bound $\lVert\theta_{j}\rVert\leq\delta$ can be considered. The considered setup allows us to model multiple independent errors affecting different qubits: For example, in a $2$-qubit system, independent Pauli-$Z$ rotation errors on each qubit can be captured by $B_{j,1}=Z\otimes I$, $B_{j,2}=I\otimes Z$. Further, we write $G_{j}(\theta_{j})$ and $G(\theta)$ to emphasize the dependence of $G_{j}$ and $G$ on the error $\theta$.
We note that the computational procedures presented in the following can be extended to more general uncertainty descriptions, following ideas from robust quantum control [37]. For example, one can also handle different bounds $\delta$ and numbers $\ell$ for each error Hamiltonian $H_{\mathrm{e},j}$, different vector norms for the bound $\lVert\theta_{j}\rVert\leq\frac{\delta}{\ell}$, or even fully uncertain Hamiltonians $H_{\mathrm{e},j}$ without any structural knowledge of the $B_{j,k}$’s. Moreover, while the main focus of this work lies on independent errors, for which $\theta_{j}$ differs for $j=1,\dots,N$, systematic errors can be analyzed as well, see Section 4.1 for an example.
@@ B.5 Discussion and comparison
It is important to emphasize that none of these approaches scale directly to high numbers of qubits or gates. In particular, each method requires to classically evaluate the matrix $G(\theta)$, whose size grows exponentially in the number of qubits. Moreover, the computation (72) additionally scales exponentially with the number of gates. In Appendix C, we address this problem via a circuit partitioning strategy, which allows us to compute bounds on $\gamma$ in a scalable fashion, paving the way for relevant problem sizes beyond classical simulation.
@@ D.1 Special case: coherent control errors
In the following, we compare the worst-case fidelity bound from Theorem 2.1 to [30], which provides a worst-case fidelity bound for the special case of coherent control errors, compare (46). The main fidelity bound from [30, Theorem 2] states
@@
Note the difference between (15) and (85). When $\delta N$ is small compared to $\gamma$, both infidelity bounds scale as $\delta^{2}N^{2}$. However, for robust algorithms with $\gamma\approx 0$, the bound (15) derived in the present paper scales in a significantly more favorable way with $\delta^{4}N^{4}$. In Section 4.2, we compare both bounds for a concrete set of algorithms, showing that the bound obtained via Theorem 2.1 is substantially less conservative than the bound by [30].